# Asymptotic structure of scalar-Maxwell theory at the null boundary

### v1

@@ Abstract
We apply the Hamiltonian formalism to investigate the massless sector of scalar field theory coupled with Maxwell electrodynamics through the Pontryagin term. Specifically, we analyze asymptotic symmetries at the null infinity of this theory, conserved charges, and their algebra. We find that the theory possesses asymptotic shift symmetries of the fields not present in the bulk manifold coming from the zero modes of the symplectic matrix of constraints. Consequently, we conclude that the real scalar field also contains asymptotic symmetries previously found in the literature by a different approach. We show that these symmetries are the origin of the electric-magnetic duality in electromagnetism with the topological Pontryagin term, and obtain non-trivial central extension between the electric and magnetic conserved charges. Finally, we examine the full interacting theory and find that, due to the interaction, the symmetry generators are more difficult to identify among the constraints, such that we obtain them in the weak-coupling limit. We find that the asymptotic structure of the theory simplifies due to a fast fall-off of the scalar field, leading to decoupled scalar and Maxwell asymptotic sectors, and losing the electric-magnetic duality.
@@ 2 Local and asymptotic symmetries in canonical formalism
where $\mathcal{H}_{0}$ does not depend explicitly on the constraints. For simplicity, we treat the spatial coordinates $x^{i}$ on the same footing as discrete indices: $(s,x^{i})\to s$. Here, $\chi_{s}\approx 0$ are all constraints in the theory, and $\lambda^{s}$ are corresponding multipliers. The symbol $\approx$’ denotes the weak equality, used for the quantities that vanish only on the constraint surface, and whose Poisson brackets (PBs) are non vanishing in general. The phase space where all constraints are solved and the canonical variables are reduced only to the independent ones, is called the reduced phase space. It is obtained when the full phase space is restricted to the constraint surface. [^5] On the constraint surface embedded in the phase space, the PBs generate evolution and local transformations. In turn, the Dirac brackets generate the evolution in the reduced phase space, where the constraints are strongly set to zero. This is a standard procedure to eliminate unphysical variables from the phase space. However, applying this standard method without any modification would lose the information about the asymptotic symmetries we are interested in, so we will keep the phase space general until we find all conserved charges.
@@
where the non-homogeneous part of the equation is computed from $Y_{s}=\left\{\mathcal{H}_{0},\chi_{s}\right\}$. The last equation shows that the nature of the multipliers depends only on the symplectic matrix. When the matrix $\Omega$ is invertible (it has maximal rank), there is a unique solution in the multipliers $\lambda=\bar{\lambda}$, and they are fully determined. If the symplectic matrix is equal to zero, then all multipliers are arbitrary, and we denoted them as $\Lambda$. In general, the number of $\Lambda$ is equal to the number of zero modes of the symplectic matrix, while a number of the solved ones, $\bar{\lambda}$, is equal to the rank of the matrix $\Omega$.
@@
As previously mentioned, all indices have a continuous part, $s\rightarrow(s,x^{i})$, where $x^{i}$ are spatial coordinates. This means that, in the calculations above, we neglected all boundary terms and the differentiability of the functional $G[\eta]$ is not guaranteed. This is resolved via the Regge-Teitelboim method [11], by adding a boundary term $Q[\eta]$ to the smeared generator,
@@ 2.1 Setup and notation
Here, $r\geq 0$ is the radial coordinate and $y^{A}$ are angles of the unit 2-sphere with the line element $\mathrm{d}\Omega^{2}=\gamma_{AB}(y)\,\mathrm{d}y^{A}\mathrm{d}y^{B}$. We also introduced the retarded time, $\epsilon=1$, and advanced time, $\epsilon=-1$, related to the usual time $t$ as $u=t-\epsilon r\in\mathbb{R}$. In that way, we have an explicit realization of the null-foliated flat spacetime, with the waves ($\mathrm{d}s=0$) propagating along the lines $u=const$ and $u+2\epsilon r=const$. The spacetime manifold has the topology $M\cong\mathbb{R}\times\Sigma$, valid at least locally, where $\Sigma$ is the spatial section parametrized by $x^{i}=(u,y^{A})$. Chosen coordinates correspond to the future cone when $\epsilon=1$, and to the past cone when $\epsilon=-1$.
@@ 3 Scalar field theory
with the shorthand notation $\phi\equiv\phi(u,r,y)$ and $\pi^{\prime}\equiv\pi(u,r^{\prime},y^{\prime})$, and the Dirac delta function $\delta^{(3)}$ is defined on the spatial section at constant $u$.
The constraint (3.2) is not present when the evolution in studied in usual timelike foliation of spacetime. Its appearance is a special feature of the time being null. Besides, the PBs between the primary constraints,
@@
We identify the matrix $Z$ in (2.5) as $1/r$. The fact that the zero mode appears at the order $1/r$ in the radial expansion suggests that the first-class constraint and the corresponding symmetry will also become relevant at the order $1/r$ of the local parameter, as suggested by (2.10).
@@
Since there exists a non-trivial zero mode of the symplectic matrix, the corresponding first-class constraint in $\chi$ generates an asymptotic symmetry. To find this constraint, we introduce a smeared generator defined on the spatial manifold according to (2.10) as
@@
when $\xi$ is field-independent ($\delta\xi=0$). The integral is defined on the celestial sphere $\mathbb{S}^{2}$ located at the boundary $r\to\infty$. Because the charge is observable, it has to be finite, as we prove below.
@@
We recognize the shift symmetry of the scalar field, the way Goldstone particles change in the vacuum. The fall-off of the parameter $\xi=\mathcal{O}(r^{-1})$, as defined by (3.13), is consistent with the above laws. Furthermore, invariance of the first Hamilton’s equation (3.11) under these transformations implies the following change of the Hamilton’s multiplier,
$\delta_{\xi}\lambda=\dot{\xi}\,,$ (3.18)
@@
$\delta_{\xi}\phi_{(1)}=\epsilon\,\xi_{(1)}\,,\qquad\delta_{\xi}\pi_{(0)}=-\sqrt{\gamma}\,\xi_{(1)}\,,\qquad\delta_{\xi}\Lambda=\dot{\xi}_{(1)}\,,$ (3.19)
@@
The conserved quantity for a radiating (massless) system is given as the total flux $\int\mathrm{d}u\,\partial_{u}Q_{K}$ [18] over the boundary. In the future cone ($\epsilon=+1$) bounded by the future null infinity $\mathscr{I}^{+}$, this integral gives $Q^{+}=Q_{K}|_{\mathscr{I}^{+}_{+}}-Q_{K}|_{\mathscr{I}^{+}_{-}}$, where $\mathscr{I}^{+}_{+}$$\mathscr{I}^{+}_{-}$ is the future/past boundary of $\mathscr{I}^{+}$. Since there are no massive particles in the spectrum, $Q_{K}|_{\mathscr{I}^{+}_{+}}=0$, and the resulting quantity coincides with the Hamiltonian charge (3.20) evaluated at the past of the future light-like infinity. The obtained charge also matches the one discussed in [7] concerning the soft theorem for the scalar field.
Similar analysis can be repeated for the past cone ($\epsilon=-1$) with the past null boundary $\mathscr{I}^{-}$, whose past $\mathscr{I}^{-}_{-}$ is radiation free, and find $Q^{-}$. From the total flux, then, we can write the usual conservation law of the massless system, $Q^{+}=Q^{-}$, provided suitable matching conditions. The antipodal matching conditions $\phi_{(1)}(u\to-\infty,y)|_{\epsilon=1}=\phi_{(1)}(u\to\infty,-y)|_{\epsilon=-1}$ satisfy continuously the asymptotic massless Klein-Gordon equation in the matching points at the null infinity, which is in agreement with [19, 12, 7].
@@
without a central extension.
The next section will show more details about the computation of these algebras.
@@ 4 Electromagnetism with the Pontryagin term
Having analyzed the pure scalar field symmetries of (1.1) in the previous section, now we can focus on the other extreme and assume the constant scalar field, $\phi=\phi_{0}=const$. This constant can be absorbed into the coupling parameter, $\theta\phi_{0}\to\theta$. We obtain the Maxwell’s action with the Pontryagin term in the background $\mathfrak{g}_{\mu\nu}$,
@@
with the local parameters $\alpha(x)$, $\alpha_{u}(x)$ and $\eta_{A}(u,y)$. While $\alpha,\alpha_{u}$ are completely arbitrary parameters that will give a usual $\mathrm{U}(1)$ gauge symmetry, the parameter $\eta$ is only asymptotic because it does not depend on the radial coordinate. The condition $\partial_{r}\eta=0$ is necessary to ensure that $S[\eta]$ is the first class variable, because the constraint $\chi^{A}$ is a combination of first and second class constraint, and the latter ones do not generate symmetries. This can be seen directly from (2.5) and (2.10), because the zero modes $v_{A}$ and the local parameters $\eta_{A}$ have the same form.
Indeed, the generators $G[\alpha]$ and $S[\eta]$ are Abelian. In particular, $S[\eta]$ commute thanks to the $r$-independence of the local parameter,
@@
Hamilton’s equations (4.14) and (4.15) are invariant under transformatins generated by $G[\alpha]$, and asymptotically invariant under transformations generated by $S[\eta]$, if the multipliers transform as
$\delta_{\alpha,\eta}\lambda_{u}=\dot{\alpha}_{u}\,,\qquad\delta_{\alpha,\eta}\lambda_{A}=\dot{\eta}_{A}\,.$ (4.29)
@@
where integrals are evaluated over $\mathbb{S}^{2}$ at infinity. For the field-independent parameters $\alpha$ and $\eta_{A}$, these charges are integrable,
@@
Because the integrals are over the sphere at infinity, the charges depend only on $\pi^{r}_{(0)}$ and $A_{A(0)}$. They are also the generators in the reduced phase space because $G_{Q}\approx Q$ and $S_{Q}\approx Q_{S}$.
@@
where $\delta$ is the exterior derivative on the symplectic manifold. Vector fields $X_{\alpha}$ and $X_{\eta}$ on this manifold are by definition
@@ 4.1 Electric-magnetic duality
defined as electric and magnetic fluxes of the electromagnetic field through the 2-sphere at the infinity. [^9] Again, in all charges, we have to take the limit $u\rightarrow-\epsilon\infty$ to obtain the total flux over the null infinity that enters the conservation law of a radiating system, as explained in Section 3 (see the text after (3.20)). The factor $r^{2}$ comes from $\sqrt{\mathfrak{g}}$ and it is essential to make the electric charge density IR finite. When $\alpha=const$, the magnetic charge becomes a topological invariant, and $Q_{\mathrm{mag}}$ reduces to a topological number.
@@
Integration in $u$ leads to [^10] Up to an arbitrary angle-dependent function that is not relevant in this computation, as it captures arbitrariness due to an asymptotic symmetry.
$\pi^{r}=\frac{\epsilon}{e^{2}}\,\sqrt{\gamma}\,\nabla^{A}A_{A}+\mathcal{O}\left(r^{-1}\right)\,.$ (4.45)
Since $\pi^{r}=\frac{r^{2}\sqrt{\gamma}}{e^{2}}\,F_{ur}$, we get
@@
recognizing the electric charge density. In sum, the Hamiltonian charges, when $\theta=0$, are
$Q^{0}[\alpha]$ $=$ $Q_{\mathrm{el}}[\alpha]\,,$
$Q_{S}^{0}[\eta]$ $=$ $Q_{\mathrm{el}}[\beta]+Q_{\mathrm{mag}}[\psi]\,,$ (4.47)
which means that $Q^{0}$ is just the standard electric charge, as expected, while $Q_{S}^{0}$ is the magnetic charge up to an additional electric term.
@@
$\left\{Q_{\mathrm{el}}[\alpha],Q_{\mathrm{mag}}[\psi]\right\}=C_{\mathrm{EM}}[\alpha,\psi]\,,$ (4.50)
and all other PBs are zero. In [17], the authors also discuss the electric-magnetic duality in electromagnetism and the corresponding charge algebra, finding the central charge $C_{\mathrm{EM}}\propto\int\mathrm{d}^{2}y\,\epsilon^{AB}\partial_{A}\psi\partial_{B}\alpha=\oint\psi\mathrm{d}\alpha$. They obtain only the $\eta(0,\psi)$-term of the central charge, which vanishes in our case, because we choose configurations with the smooth boundary of the 2-sphere, leaving only the $\eta(\beta,0)$-term. In contrast, in [17], the authors look at non-trivial configurations with non-smooth boundaries, such as the point-like electric charge that plays the role of a monopole for a dual potential.
@@
which in the asymptotic expansion leads to
@@
such that, replacing $\pi^{r}$ from the Hamilton’s equation, we find that the $\theta$-term does not affect the identity (4.46). Hence, we get
$Q[\alpha]$ $=$ $Q_{\mathrm{el}}[\alpha]-Q_{\mathrm{mag}}[\theta\alpha]\,,$
$Q_{S}[\eta]$ $=$ $Q_{\mathrm{el}}[\beta-\theta\,\psi]+Q_{\mathrm{mag}}[\psi+\theta\beta]\,.$ (4.53)
@@
$Q_{\mathrm{tot}}[\xi]=Q[\alpha]+Q_{S}[\eta]=\oint\limits_{\mathbb{S}_{\infty}^{2}}\mathrm{d}^{2}y\,\sqrt{\gamma}\,\xi^{T}q\,,$ (4.54)
@@
where we neglected the boundary terms taken over the 2-sphere.
To this end, we have to invert eqs. (4.53) and express the electric and magnetic charges in terms of $Q$ and $Q_{S}$, which is ambiguous, since there are more parameters than charges. One parameter can always be set to zero, except $\alpha$, otherwise the relations are not invertible. It means the algebra can look differently in different bases of charges $(Q,Q_{S})$, but the total charge (4.54) remains unchanged.
To choose which parameter to set to zero, we remember from the $\theta=0$ case given by (4.47) that $\alpha$ and $\beta$ have an electric nature and $\psi$ magnetic nature. The introduction of $\theta\neq 0$ mixes them up and they all contribute to both electric and magnetic terms, but it is more natural to set $\beta=0$ to be able to study $\theta\rightarrow 0$ limit at the end. Then, from (4.55), we can express $\alpha$ and $\psi$ in terms of the true electric and magnetic gauge parameters $\xi_{\mathrm{el}}$ and $\xi_{\mathrm{mag}}$ as

### v2

@@ Abstract
We apply the Hamiltonian formalism to investigate the massless sector of scalar field theory coupled with Maxwell electrodynamics through the Pontryagin term. To this end, we generalize the Dirac procedure to include radially independent zero modes of the symplectic matrix associated with the asymptotic symmetries. Specifically, we analyze asymptotic symmetries at the null infinity of this theory, conserved charges, and their algebra. We find that the theory possesses asymptotic shift symmetries of the fields not present in the bulk manifold coming from the zero modes of the symplectic matrix of constraints. Consequently, we conclude that the real scalar field also contains asymptotic symmetries previously found in the literature by a different approach. We show that these symmetries can be seen as the electric-magnetic duality in electromagnetism with the topological Pontryagin term, and obtain a non-trivial central extension between the electric and magnetic conserved charges. Finally, we examine the full interacting theory and find that, due to the interaction, the symmetry generators are more difficult to identify among the constraints, such that we obtain them in the weak-coupling limit. We find that the asymptotic structure of the theory simplifies due to a fast fall-off of the scalar field, leading to decoupled scalar and Maxwell asymptotic sectors, and losing the electric-magnetic duality.
@@ 2 Local and asymptotic symmetries in canonical formalism
where $\mathcal{H}_{0}$ does not depend explicitly on the constraints. For simplicity, we treat the coordinates $x^{i}$ that exclude $u$ on the same footing as discrete indices: $(s,x^{i})\to s$. Here, $\chi_{s}\approx 0$ are all constraints in the theory, and $\lambda^{s}$ are corresponding multipliers. The symbol $\approx$’ denotes the weak equality, used for the quantities that vanish only on the constraint surface, and whose Poisson brackets (PBs) are non vanishing in general. The phase space where all constraints are solved and the canonical variables are reduced only to the independent ones, is called the reduced phase space. It is obtained when the full phase space is restricted to the constraint surface. [^5] On the constraint surface embedded in the phase space, the PBs generate evolution and local transformations. In turn, the Dirac brackets generate the evolution in the reduced phase space, where the constraints are strongly set to zero. This is a standard procedure to eliminate unphysical variables from the phase space. However, applying this standard method without any modification would lose the information about the asymptotic symmetries we are interested in, so we will keep the phase space general until we find all conserved charges.
@@
where the non-homogeneous part of the equation is computed from $Y_{s}=\left\{\mathcal{H}_{0},\chi_{s}\right\}$. The last equation shows that the nature of the multipliers depends only on the symplectic matrix. When the matrix $\Omega$ is invertible (it has maximal rank), there is a unique solution in the multipliers $\lambda=\bar{\lambda}$, and they are fully determined. If the symplectic matrix is equal to zero, then all multipliers are arbitrary, and we denoted them as $\Lambda$. In the finite-dimensional case, the dimension of the kernel of $\Omega_{ss^{\prime}}$ specifies the number of independent functions $\Lambda$, while the rank of $\Omega_{ss^{\prime}}$ determines the number of independent functions $\bar{\lambda}$. However, when the matrix $\Omega_{ss^{\prime}}(x,x^{\prime})$ is infinite-dimensional, these numbers are more difficult to count, so the appropriate characterization is given by the dimension of the kernel, the rank of $\Omega$ and the specification of a (sub)manifold where they are defined.
@@
As previously mentioned, all indices have a continuous part, $s\rightarrow(s,x^{i})$, where $x^{i}$ are all coordinates except the retarded/advanced time. This means that, in the calculations above, we neglected all boundary terms and the differentiability of the functional $G[\eta]$ is not guaranteed. This is resolved via the Regge-Teitelboim method [11], by adding a boundary term $Q[\eta]$ to the smeared generator,
@@ 2.1 Setup and notation
Here, $r\geq 0$ is the radial coordinate and $y^{A}$ are angles of the unit 2-sphere with the line element $\mathrm{d}\Omega^{2}=\gamma_{AB}(y)\,\mathrm{d}y^{A}\mathrm{d}y^{B}$. We also introduced the retarded time, $\epsilon=1$, and advanced time, $\epsilon=-1$, related to the usual time $t$ as $u=t-\epsilon r\in\mathbb{R}$. In that way, we have an explicit realization of the null-foliated flat spacetime, with the waves ($\mathrm{d}s=0$) propagating along the lines $u=\mathrm{const.}$ and $u+2\epsilon r=\mathrm{const}$. The spacetime manifold has the topology $M\cong\mathbb{R}\times\Sigma$, valid at least locally, where $\Sigma$ is the null surface parametrized by  [^6] The null coordinates $\ell_{\mu}\mathrm{d}x^{\mu}=\mathrm{d}u$ and $n_{\mu}\mathrm{d}x^{\mu}=\mathrm{d}u+2\epsilon\,\mathrm{d}r$ define the null vectors $\ell_{\mu}=\partial_{\mu}u=(1,0,0,0)$ and $n_{\mu}=(1,2\epsilon,0,0)$. Thus, the $u=\mathrm{const}.$ surface is lightlike, i.e., with the normal vector $\ell_{\mu}$ satisfying $g^{\mu\nu}\ell_{\mu}\ell_{\nu}=0$. $x^{i}=(r,y^{A})$. Chosen coordinates correspond to the future cone when $\epsilon=1$, and to the past cone when $\epsilon=-1$.
@@ 3 Scalar field theory
with the shorthand notation $\phi\equiv\phi(u,r,y)$ and $\pi^{\prime}\equiv\pi(u,r^{\prime},y^{\prime})$, and the Dirac delta function $\delta^{(3)}$ is defined on the null surface with constant $u$ in spherical coordinates. The constraint (3.2) is not present when the evolution is studied in the usual timelike foliation of spacetime. Its appearance is a special feature of the time being null. Besides, the PBs between the primary constraints,
@@
Therefore, the kernel of the matrix $\Omega$ has dimension one and its rank is zero in the subspace $r=\mathrm{const}.$, while its rank is one in the full spacetime. We identify the matrix $Z$ in (2.5) as $1/r$. The fact that the zero mode appears at the order $1/r$ in the radial expansion suggests that the first-class constraint and the corresponding symmetry will also become relevant at the order $1/r$ of the local parameter, as suggested by (2.10).
@@
Since there exists a non-trivial zero mode of the symplectic matrix, the corresponding first-class constraint in $\chi$ generates an asymptotic symmetry. To find this constraint, we introduce a smeared generator defined on the null manifold according to (2.10) as
@@
when $\xi$ is field-independent ($\delta\xi=0$). The integral is defined on the celestial sphere $\mathbb{S}^{2}$ located at the boundary $r\to\infty$. As regards the time dependence, it is specified as follows.
For a radiating massless system, the charge is defined as the total flux $\int\mathrm{d}u\,\partial_{u}Q_{K}$ [17] over the boundary that, in the future cone ($\epsilon=1$), corresponds to $-Q_{K}|_{\mathscr{I}_{-}^{+}}$ in absence of massive particles in the spectrum ($Q_{K}|_{\mathscr{I}_{+}^{+}}=0$). This means that the charge (3.16) cannot be evaluated at arbitrary time $u=u_{0}=\mathrm{const}$, but at $u\to-\infty$. This is secured if one completes the integral (3.13) by adding a patch of null infinity from $u_{0}$ to $u\rightarrow-\infty$. The final charge $Q^{+}$ defined only over the boundary of the future null infinity is
$Q^{+}[\xi]=-r^{2}\oint_{\mathscr{I}_{-}^{+}}\mathrm{d}^{2}y\,\sqrt{\gamma}\,\xi\phi\,.$ (3.17)
Doing so, we also ensure that the integral is always performed over a Cauchy surface (see [18, 19, 20, 21, 22] for a deeper discussion). In practice, it means that all boundary quantities in the charge are $u$-independent (evaluated at $u\rightarrow-\infty$). Because $\xi_{(1)}$ does not depend on $r$, it implies that it is $u$-independent in the whole spacetime, namely, $\xi=\frac{\xi_{(1)}(y)}{r}$. The charge $Q^{-}$ similar to (3.17) can be defined in the past cone ($\epsilon=-1$), with the global sign positive and the integral evaluated at $\mathscr{I}_{+}^{-}$, that is, in the limit $u\to+\infty$. In sum, $Q^{\pm}[\xi]=-\epsilon\,Q_{K}[\xi(y)]\,|_{u\to-\epsilon\infty}$.
To simplify the notation for all $\epsilon=\pm 1$, we will work with the usual Hamiltonian charge (3.16), assuming the $u$-independence of the asymptotic symmetry parameter and with the boundary integral evaluated at $\mathscr{I}_{+}^{-}/\mathscr{I}_{-}^{+}$. Differences in signs can always be absorbed in the local parameter.
@@
We recognize the shift symmetry of the scalar field, the way Goldstone particles change in the vacuum. The fall-off of the parameter $\xi=\mathcal{O}(r^{-1})$, as defined by (3.13), is consistent with the above laws. For instance, the invariance of Hamilton’s equation for the momentum (3.11) is satisfied only in the asymptotic limit, leading to $\dot{\xi}\rightarrow 0$ when $r\rightarrow\infty$, which is consistent with the $u$-independence of $\xi$. Furthermore, invariance of the first Hamilton’s equation (3.11) under these transformations and the fact that the parameter is $u$-independent, imply invariance of the Hamilton’s multiplier under symmetry transformations,
$\delta_{\xi}\lambda=\dot{\xi}=0\,,$ (3.19)
@@
$\delta_{\xi}\phi_{(1)}=\epsilon\,\xi_{(1)}\,,\qquad\delta_{\xi}\pi_{(0)}=-\sqrt{\gamma}\,\xi_{(1)}\,,\qquad\delta_{\xi}\Lambda=0\,,$ (3.20)
@@
As discussed previously, the conserved quantity for a radiating system in the future cone coincides with the Hamiltonian charge (3.21) evaluated at the past of the future light-like infinity, $Q^{+}=-Q_{K}|_{\mathscr{I}^{+}_{-}}$. This charge also matches the one discussed in [7] concerning the soft theorem for the scalar field. Similarly, the conserved quantity for a radiating system in the past cone is the Hamiltonian charge evaluated at the future of the past light-like infinity, $Q^{-}=Q_{K}|_{\mathscr{I}^{-}_{+}}$. From the total flux, then, we can write the usual conservation law of the massless system, $Q^{+}=Q^{-}$, provided suitable matching conditions. The antipodal matching conditions $\phi_{(1)}(u\to-\infty,y)|_{\epsilon=1}=\phi_{(1)}(u\to\infty,-y)|_{\epsilon=-1}$ and $\xi(y)|_{\epsilon=1}=-\xi(-y)|_{\epsilon=-1}$ (the last one due to our definition of the parameter given by eq. (3.18)) satisfy continuously the asymptotic massless Klein-Gordon equation in the matching points at the null infinity, which is in agreement with [23, 12, 7].
@@
without a central extension. The next section will show more details about the computation of these algebras.
@@ 4 Electromagnetism with the Pontryagin term
Having analyzed the pure scalar field symmetries of (1.1) in the previous section, now we can focus on the other extreme and assume the constant scalar field, $\phi=\phi_{0}=\mathrm{const}$. This constant can be absorbed into the coupling parameter, $\theta\phi_{0}\to\theta$. We obtain the Maxwell’s action with the Pontryagin term in the background $\mathfrak{g}_{\mu\nu}$,
@@
with the local parameters $\alpha(x)$, $\alpha_{u}(x)$ and $\eta_{A}(y)$. The above integrals defined in $u=u_{0}$ are completed by adding corresponding patches of null infinity from the point $u=u_{0}$ to $u\rightarrow-\epsilon\infty$.
While $\alpha,\alpha_{u}$ are completely arbitrary parameters that will give a usual $\mathrm{U}(1)$ gauge symmetry, the parameter $\eta$ is only asymptotic because it does not depend on the radial coordinate. The condition $\partial_{r}\eta=0$ is necessary to ensure that $S[\eta]$ is the first class variable, because the constraint $\chi^{A}$ is a combination of first and second class constraint, and the latter ones do not generate symmetries. This can be seen directly from (2.5) and (2.10), because the zero modes $v_{A}$ and the local parameters $\eta_{A}$ have the same form. Furthermore, $\eta_{A}$ is also $u$-independent. Namely, as we saw in Sec. 3, this condition is satisfied for the quantities defined at the null boundary, and for $\eta_{A}$ also in the bulk because it is $r$-independent.
Indeed, the generators $G[\alpha]$ and $S[\eta]$ are first class. In particular, $S[\eta]$ commute thanks to the $r$-independence of the local parameter,
@@
Hamilton’s equations (4.14) and (4.15) are invariant under transformations generated by $G[\alpha]$, and asymptotically invariant under transformations generated by $S[\eta]$, if the multipliers transform as
$\delta_{\alpha,\eta}\lambda_{u}=\dot{\alpha}_{u}\,,\qquad\delta_{\alpha,\eta}\lambda_{A}=\dot{\eta}_{A}=0\,.$ (4.29)
@@
where integrals are evaluated over $\mathbb{S}^{2}$ at the null infinity $\mathscr{I}^{+}_{-}$ for the future cone and $\mathscr{I}^{-}_{+}$ for the past cone. For the field-independent parameters $\alpha$ and $\eta_{A}$, these charges are integrable,
@@
Because the integrals have to be taken in the limits $r\to\infty$ and $u\to-\epsilon\infty$, the charges depend only on $\pi^{r}_{(0)}(y)$ and $A_{A(0)}(y)$. They are also the generators in the reduced phase space because $G_{Q}\approx Q$ and $S_{Q}\approx Q_{S}$.
@@
where $\delta$ is the exterior derivative on the symplectic manifold. Similarly as previously discussed for smeared generators, the symplectic form (4.34) has to be integrated over a Cauchy surface, such that the integral has to include a patch of null infinity from $u=\mathrm{const}.$ to $u\to\mp\infty$ for the future/past cone. In this way, the obtained charges are indeed evaluated at $\mathscr{I}^{+}_{-}/\mathscr{I}^{-}_{+}$, as argued. We will not write this patch explicitly. For details, see [18, 19, 20, 21, 22].  [^10] Strictly speaking, (4.34) is a presymplectic form because it is degenerate due to isometries generated by (4.36) with the parameter $\eta_{A}$ that has no support at $\mathscr{I}^{+}_{-}/\mathscr{I}^{-}_{+}$. Therefore, they have to be quotient out from the covariant configuration space [27] by requiring that all the parameters are $u$-independent on the null boundary. Namely, it can be shown that the quotient space is the space of time-independent functions. We thank to the anonymous referee for clarifying this issue to us.
Vector fields $X_{\alpha}$ and $X_{\eta}$ on this manifold are by definition
@@ 4.1 Electric-magnetic duality
defined as electric and magnetic fluxes of the electromagnetic field through the 2-sphere at the infinity. [^11] Again, in all charges, we have to take the limit $u\rightarrow-\epsilon\infty$ to obtain the total flux over the null infinity that enters the conservation law of a radiating system, as explained in Section 3 (see the text after (3.21)). The factor $r^{2}$ comes from $\sqrt{\mathfrak{g}}$ and it is essential to make the electric charge density IR finite. When $\alpha=\mathrm{const.}$, the magnetic charge becomes a topological invariant, and $Q_{\mathrm{mag}}$ reduces to a topological number.
@@
Integration in $u$ leads to
$\pi^{r}=\frac{\epsilon}{e^{2}}\,\sqrt{\gamma}\,\nabla^{A}A_{A}+k(y)+\mathcal{O}\left(r^{-1}\right)\,$ (4.45)
up to an arbitrary angle-dependent function which we set to zero for simplicity, $k(y)=0$. Since $\pi^{r}=\frac{r^{2}\sqrt{\gamma}}{e^{2}}\,F_{ur}$, we get
@@
recognizing the electric charge density. In sum, the Hamiltonian charges, when $\theta=0$, are identified with the Lagrangian quantities through the mapping
$Q^{0}[\alpha]$$\equiv$ $Q_{\mathrm{el}}[\alpha]\,,$
$Q_{S}^{0}[\eta]$$\equiv$ $Q_{\mathrm{el}}[\beta]+Q_{\mathrm{mag}}[\psi]\,,$ (4.47)
which means that $Q^{0}$ is just the standard electric charge, while $Q_{S}^{0}$ is the magnetic charge up to an additional electric term.
@@
$\left\{Q_{\mathrm{el}}[\alpha],Q_{\mathrm{mag}}[\psi]\right\}^{*}=C_{\mathrm{EM}}[\alpha,\psi]\,,$ (4.50)
and all other PBs are zero.
It is worth emphasizing that the mapping (4.47) that identifies the electric and magnetic fluxes with the Hamiltonian charges is not invertible when $\alpha=0$ or $\psi=0$, producing ambiguities in the definition of charges. For instance, we could have $Q_{\mathrm{el}}[\alpha]=Q^{0}[\alpha]$ or $Q^{\prime}_{\mathrm{el}}[\alpha]=Q_{S}^{0}[\eta(\alpha,\psi=0)]$ for the electric flux. The corresponding Abelian algebras (4.48) would have different central extensions. Thus, we have to require $\alpha,\psi\neq 0$ to have invertible mapping $(Q_{\mathrm{el}},Q_{\mathrm{mag}})\leftrightarrow(Q^{0},Q^{0}_{S})$. Then the $\mathrm{U}(1)$ charge is given by a usual expression in both Hamiltonian and Lagrangian formalisms.
In [21, 29], the authors also discuss the electric-magnetic duality in electromagnetism and the corresponding charge algebra, finding the central charge $C_{\mathrm{EM}}\propto\int\mathrm{d}^{2}y\,\epsilon^{AB}\partial_{A}\psi\partial_{B}\alpha=\oint\psi\mathrm{d}\alpha$. They obtain only the $\eta(0,\psi)$-term of the central charge, which vanishes in our case, because we choose configurations with the smooth boundary of the 2-sphere, leaving only the $\eta(\beta,0)$-term. In contrast, in [29], the authors look at non-trivial configurations with non-smooth boundaries, such as the point-like electric charge that plays the role of a monopole for a dual potential. The electric-magnetic duality in the context of soft charges has also been discussed in [21].
@@
which in the asymptotic expansion leads to [^12] Similarly as in (4.45), we neglect the integration constant, $k(y)=0$.
@@
such that, replacing $\pi^{r}$ from the Hamilton’s equation, we find that the $\theta$-term does not affect the identity (4.46). Hence, we get the identifications that map Lagrangian to Hamiltonian charges,
$Q[\alpha]$$\equiv$ $Q_{\mathrm{el}}[\alpha]-Q_{\mathrm{mag}}[\theta\alpha]\,,$
$Q_{S}[\eta]$$\equiv$ $Q_{\mathrm{el}}[\beta-\theta\,\psi]+Q_{\mathrm{mag}}[\psi+\theta\beta]\,.$ (4.53)
@@
$Q_{\mathrm{tot}}[\xi]=Q[\alpha]+Q_{S}[\eta]\equiv\oint\limits_{\mathbb{S}_{\infty}^{2}}\mathrm{d}^{2}y\,\sqrt{\gamma}\,\xi^{T}q\,,$ (4.54)
@@
where we neglected the boundary terms taken over the 2-sphere at the null infinity.
To this end, we have to invert eqs. (4.53) and express the electric and magnetic charges in terms of $Q$ and $Q_{S}$, which is ambiguous, since there are more parameters than charges. Thanks to $\mathrm{SL}(2,\mathbb{R})$ symmetry, one parameter can always be set to zero, except $\alpha$, otherwise the relations (4.53) are not invertible. To choose which parameter to set to zero, we remember from the $\theta=0$ case given by (4.47) that $\alpha$ and $\beta$ have an electric nature and $\psi$ magnetic nature. The introduction of $\theta\neq 0$ mixes them up and they all contribute to both electric and magnetic terms, but it is more natural to set $\beta=0$ to be able to study $\theta\rightarrow 0$ limit at the end. Then, from (4.55), we can express $\alpha$ and $\psi$ in terms of the true electric and magnetic gauge parameters $\xi_{\mathrm{el}}$ and $\xi_{\mathrm{mag}}$ as